# Cross-lingual Transfer of Reward Models in Multilingual Alignment

### v1

@@ Abstract
Reinforcement learning with human feedback (RLHF) is shown to largely benefit from precise reward models (RMs). However, recent studies in reward modeling schemes are skewed towards English, limiting the applicability of RLHF in multilingual alignments. In this work, we investigate the cross-lingual transfer of RMs trained in diverse languages, primarily from English. Our experimental results demonstrate the strong cross-lingual transfer of English RMs, exceeding target language RMs by 3$\sim$4average increase in Multilingual RewardBench. Furthermore, we analyze the cross-lingual transfer of RMs through the representation shifts. Finally, we perform multilingual alignment to exemplify how cross-lingual transfer in RM propagates to enhanced multilingual instruction-following capability, along with extensive analyses on off-the-shelf RMs. We release the code [^1] Code - IQ-KAIST/rm-lingual-transfer, model and data [^2] Data & Models - HF Collection.
@@ 3.1 English preserves general representations
with $S$ as a singular values of $\mathcal{H}_{\theta}(x)$. Intuitively, having $f_{\theta}(x)$ close to 1 indicates the hidden states in different languages are homogeneous: i.e., representations are embedded into similar space.
@@ 4.1 Experimental Details
Figure 3: Multilingual AlpacaEval results of Qwen2.5-7B-Instruct models fine-tuned with DPO on on-policy generations for four non-English languages. The alignment data were labeled with either English RM or target language RM.
@@ 4.2 Results and Analysis
As shown in Figure 3, models aligned with English RM show a notable leap compared to Qwen2.5-7B-Instruct ("Base"), by increasing up to 11.8% point. As the win rate was measured against GPT-4-Turbo, a strong proprietary language model, such enhancements strongly support the validity of using English RMs for multilingual alignment in desired languages.
@@
We emphasize that using high-quality English preference data of better accessibility is a decent choice, considering the efficiency and efficacy in real-world cases. In Figure 3, models aligned with English RM were comparable to or exceeded ones with target language RMs, having at most 1.5% disparity in the worst scenario. Thus, adopting an English RM for multilingual alignment is a cost-efficient yet performant alternative, discarding the need for scaled translations.

### v2

@@ Abstract
Reinforcement learning with human feedback (RLHF) is shown to largely benefit from precise reward models (RMs). However, recent studies in reward modeling schemes are skewed towards English, limiting the applicability of RLHF in multilingual alignments. In this work, we investigate the cross-lingual transfer of RMs trained in diverse languages, primarily from English. Our experimental results demonstrate the strong cross-lingual transfer of English RMs, exceeding target language RMs by 3$\sim$4average increase in Mutlilingual RewardBench. Furthermore, we analyze the cross-lingual transfer of RMs through the representation shifts. Finally, we perform multilingual alignment to exemplify how cross-lingual transfer in RM propagates to enhanced multilingual instruction-following capability, along with extensive analyses on off-the-shelf RMs. We release the code [^1] Code - IQ-KAIST/rm-lingual-transfer, model and data [^2] Data & Models - HF Collection.
@@ 3.1 English preserves general representations
with $S$ as singular value matrix of $\mathcal{H}_{\theta}(x)$. Intuitively, having $f_{\theta}(x)$ close to 1 imples the hidden states in different languages are homogeneous: i.e., representations are embedded into similar space.
@@ 4.1 Experimental Details
Figure 3: Multilingual AlpacaEval results of Qwen2.5-7B-Instruct models fine-tuned with DPO on on-policy generations for four non-English languages over fine runs. The alignment data were labeled with either English RM or target language RM. Results are averaged over 5 runs.
@@ 4.2 Results and Analysis
As shown in Figure 3, models aligned with English RM show a notable leap compared to Qwen2.5-7B-Instruct ("Base"), by increasing up to 9.3% point in Spanish. As the win rate was measured against GPT-4-Turbo, a strong proprietary language model, such enhancements strongly support the validity of using English RMs for multilingual alignment in desired languages.
@@
We emphasize that using high-quality English preference data of better accessibility is a decent choice, considering the efficiency and efficacy in real-world cases. In Figure 3, models aligned with English RM outperformed or at least on par with ones with target language RMs, tied only in Chinese. Thus, adopting an English RM for multilingual alignment is a cost-efficient yet performant alternative, discarding the need for scaled translations for the reward model.